# Characterizing graph-nonedge pairs with single interval Cayley configuration spaces in 3-dimensions

### v1

@@ B.4. Proof of Proposition 4 [Chained winged-graph minors]
The lemma is immediate if $xy$ is an edge of $G\cup f$. Otherwise, assume to the contrary that $f$ is preserved in some forbidden minor $[G\cup\{f,xy\}]$. Let $[G\cup f]_{A}$ be an $f$-winged graph minor such that $xy$ is one of its pins. By Lemma 12, we can assume wlog that $[G\cup f]_{A}$ has wing tips $s$ and $t$. Since $[G\cup f]_{A}$ has at least two pins, by Proposition 3, $C=\{x,y,t\}$ is a separator of $G\cup\{f,xy\}$. Furthermore, since every forbidden minor is $4$-connected and has at least five vertices, there is a unique $C$-component $H$ of $G\cup\{f,xy\}$ that contains some vertex $w$ such that $[w]$ is not contained in $[C]$. If $H$ contains $f$, then the paths in $G\cup f$ given by the existence of $[G\cup f]_{A}$ imply the existence of some minor $[G\cup f]_{B}$ in which $H$ is preserved and $[C]_{B}$ is a clique. Combining these facts shows that $[G\cup f]_{B}$ has an $[f]_{B}$-preserving forbidden minor, and thus $G\cup f$ has an $f$-preserving forbidden minor, which is a contradiction.
Next, if $H$ does not contain $f$, then the existence of $[G\cup\{f,xy\}]$ implies the existence of a minor $[G\cup\{f,xy\}]_{D}$ that preserves $H$, has an $[f]_{D}$-preserving forbidden minor, and has the property that $[w]_{D}$ is contained in $[C]_{D}$ for every vertex $w\in V(G)\setminus V(H)$. Hence, both endpoints of $[f]_{D}$ are contained $[C]_{D}$. If $[f]_{D}=[xy]_{D}$, then $xy$ is doubled in $[G\cup\{f,xy\}]_{D}$. Otherwise, assume wlog that $[f]_{D}=[xt]_{D}$. Then, we must have $[s]_{D}=[x]_{D}$. Since $\{x,y,t\}$ is a separator of $G\cup f$, $[y]_{A}^{-1}\setminus V(H)$ must be a subset of $[x]_{D}^{-1}\cup[y]_{D}^{-1}$. Therefore, it is easy to see that $xy$ is doubled in $[G\cup\{f,xy\}]_{D}$. Thus, $xy$ is doubled in $[G\cup\{f,xy\}]_{D}$ in either case, and so $G\cup f$ clearly has an $f$-preserving forbidden minor, which is a contradiction. Thus, $G\cup\{f,xy\}$ has no $f$-preserving forbidden minor. ∎
@@
Consider any winged graph pair $(G,f)$ and any pin $xy$ of an $f$-winged graph minor $[G\cup f]$. If the converse direction of Theorem 4 is true for any graph-nonedge pair whose graph has strictly fewer vertices than $G$, then $(G\setminus[f]^{-1},xy)$ has the $3$-SIP.
@@
The lemma is immediate if $xy$ is an edge of $G\setminus[f]^{-1}$. Otherwise, note that $\{x,y\}$ is a clique separator of $(G\setminus[f]^{-1})\cup xy$. Hence, by Lemma 3, it suffices to show that show that $(H\setminus xy,xy)$ has the $3$-SIP for each $\{x,y\}$-component $H$ of $(G\setminus[f]^{-1})\cup xy$ that contains $xy$. If $(H\setminus xy,xy)$ does not have the $3$-SIP, then, since $|V(H)|<|V(G)|$, the contrapositive of the converse direction of Theorem 4 shows that $H$ has an $xy$-preserving forbidden minor. However, this fact along with the paths in $G\setminus[f]^{-1}$ given by the existence of $[G\cup f]$ imply that $G\cup f$ has an $f$-preserving forbidden minor, which is a contradiction. Thus, $(H\setminus xy,xy)$ has the $3$-SIP, and the proof is complete. ∎
Proof of Proposition 4.
Let $(G^{\prime},f)$ be any $f$-winged graph pair with an $f$-winged graph minor $[G^{\prime}\cup f]$ such that (i) $G$ is a spanning subgraph of $G^{\prime}$ and $(G,f)$ has the $3$-SIP if $(G^{\prime}\setminus[f]^{-1},[f]^{-1})$ is $3$-Cayley-connected and (ii) the number $k$ of pins of $[G^{\prime}\cup f]$ that are edges of $G^{\prime}\cup f$ is maximum over all winged graph pairs satisfying (i). Assume that some pin $xy$ of $[G^{\prime}\cup f]$ is a nonedge of $G^{\prime}\cup f$. Using Lemma 16, we see that $(G^{\prime}\cup xy,f)$ is a winged graph pair. Also, the minor $[G^{\prime}\cup\{f,xy\}]_{A}$ such that $[u]_{A}=[v]_{A}$ if and only if $[u]=[v]$, for any two vertices $u,v\in V(G^{\prime})$, is an $f$-winged graph minor. Note that $G$ is a spanning subgraph of $G^{\prime}\cup xy$ and the number of pins of $[G^{\prime}\cup\{f,xy\}]_{A}$ that are edges of $G^{\prime}\cup\{f,xy\}$ is strictly greater than $k$. Also, since $G$ is a spanning subgraph of $G^{\prime}$, Lemma 17 shows that $(G^{\prime}\setminus[f]^{-1},xy)$ has the $3$-SIP. Therefore, Lemma 2 shows that $(G^{\prime}\setminus[f]^{-1},[f]^{-1})$ is $3$-Cayley connected if $((G^{\prime}\setminus[f]^{-1})\cup xy,[f]^{-1})$ is $3$-Cayley connected. Consequently, $(G,f)$ has the $3$-SIP if $((G^{\prime}\setminus[f]^{-1})\cup xy,[f]^{-1})$ is $3$-Cayley connected. Combining the above facts contradicts (ii). Thus, every pin of $[G^{\prime}\cup f]$ is an edge of $G^{\prime}\cup f$, and the proof is complete. ∎
B.5. Proof of Proposition 5[$(G,\{ux_{2},uy_{2}\})$ is $3$-Cayley-connected]
The proof of Proposition 5 requires Lemmas 19 and 22, below. Lemma 19 is proved using Lemma 18, also below.
Lemma 18 (No forbidden minor preserves endpoints of pins or $f$).
@@ B.5. Proof of Proposition 5[$(G,\{ux_{2},uy_{2}\})$ is $3$-Cayley-connected]
Assume to the contrary that $xw$ is preserved in some forbidden minor of $G\setminus[f]^{-1}$. Let $xy$ be a pin of $[G\cup f]$. Since $xy$ is an edge of $G$ and no forbidden minor has a separator of size two, some $\{x,y\}$-component of $G\setminus[f]^{-1}$ has an $xw$-preserving forbidden minor. However, combining this with the existence of $[G\cup f]$ implies that $G\cup f$ has some $xw$-preserving forbidden minor that is also $f$-preserving, which is a contradiction. Thus, the lemma is proved. ∎
Consider a winged graph pair $(G,f)$ with $f=uv$ and a chained $f$-winged graph minor $[G\cup f]$ whose ordered set of pins is $\{x_{i}y_{i}\}$. Also, let $H_{u}(x_{i}y_{i})$ and $H_{v}(x_{i}y_{i})$ be the $\{x,y\}$-components of $G\setminus[f]^{-1}$ that contain $u$ and $v$, respectively. The graphs $H_{u}(x_{1}y_{1})$ and $H_{v}(x_{n}y_{n})$ and the $\{x_{i+1},y_{i+1}\}$-component of $H_{v}(x_{i}y_{i})$ that contains $x_{i}y_{i}$ are the links of $[G\cup f]$.
@@
Proof.
Assume that some link $L$ of $[G\cup f]$ is not $3$-connected. We will show that $(G,f)$ has the $3$-SIP. First, we show that $L$ is $2$-connected. If this is not the case, then $L$ has some separator $\{w\}$. Since each pin of $[G\cup f]$ is an edge of $G$ and $[G\cup f]$ has singleton wing tips, $\{w\}$ is clearly a separator of $G$. Also, since $G\cup f$ is $2$-connected, there are exactly two $\{w\}$-components of $G$: the unique ones containing the endpoints $u$ and $v$ of $f$, respectively. Hence, we have $w\notin\{u,v\}$. Letting the vertex set of $[G\cup f]$ be $\{[a_{i}]\}$, the above facts above show that $w\notin[a_{1}]^{-1}\cup[a_{2}]^{-1}$. However, the paths in $G$ given by the existence of $[G\cup f]$ show that $u$ and $v$ are contained in the same $\{w\}$-component of $G$, which is a contradiction. Therefore, $L$ is $2$-connected.
Property (i): $\{x,y\}$ is a separator of $G$. This follows easily from the fact that each pin of $[G\cup f]$ is an edge of $G$.
Property (ii): $w$, $z$, and $w^{\prime}$ are all contained in the same $\{x,y\}$-component of $G$. Since each pin of $[G\cup f]$ is an edge of $G$, $w$ and $z$ are contained in the same $\{x,y\}$-component $H$ of $G$. Hence, if $w^{\prime}$ is not contained in $H$, then $\{x,y\}$ must be a $uv$-separator of $G$, since $\{w,z\}$ is a $uv$-separator of $G$. Along with the fact that $[G\cup f]$ has singleton wing tips and the paths in $G$ given by the existence of $[G\cup f]$, we get that $xy$ is a pin of $[G\cup f]$, which contradicts the above observation. Therefore, $w^{\prime}$ is contained in $H$.
Assume to the contrary that some $\{x,y\}$-component $H$ of $L\cup xy$ has an $xy$-preserving forbidden minor. Using Lemma 18, we see that no $\{x,y\}$-component of $L\cup xy$ has a $wz$-preserving forbidden minor. Also, $H_{wz}$ is $2$-connected since $L$ is $2$-connected. These facts easily imply that $H$ is $H_{wz}$. Proposition 1 shows that $(H_{wz}\setminus wz,wz)$ is a winged graph pair. Using Lemma 10, we see that there exists a $wz$-winged graph minor $[H_{wz}]_{A}$ such that $x$ is not contained in $[x^{\prime}]_{A}^{-1}\cup[y^{\prime}]_{A}^{-1}$. Additionally, Lemma 12 allows us to assume that $[H_{wz}]_{A}$ has wing tips $x^{\prime}$ and $y^{\prime}$. Using Lemma 18, we see that $w$, $z$, and $w^{\prime}$ must be contained in $[x^{\prime}]_{A}^{-1}\cup[y^{\prime}]_{A}^{-1}$. Consequently, $\{x^{\prime},y^{\prime}\}$ is a separator of $H_{wz}$, and hence of $L$. Wlog, assume that $[x^{\prime}x]_{A}$ is an edge. Since $xy$ is an edge of $H_{wz}$, $y$ is either contained in $[x^{\prime}]_{A}^{-1}$ or not contained in $[x^{\prime}]_{A}^{-1}\cup[y^{\prime}]_{A}^{-1}$. In the former case, it must be that $y=x^{\prime}$. Therefore, in either case, all paths from any vertex in $\{w,z,w^{\prime}\}$ to any vertex in $\{x,y\}$ contain some vertex in $\{x^{\prime},y^{\prime}\}$. However, this implies that the the $\{x^{\prime},y^{\prime}\}$-component of $L$ that contains $w$, $z$, and $w^{\prime}$ has strictly more vertices than $H_{wz}$, which contradicts our maximality assumption on $\{x,y\}$. Thus, $H_{wz}$ has no $xy$-preserving forbidden minor, and so the proof is complete. ∎
Next, Lemma 22 is proved using Lemma 21, below, and Theorems 5 and 6, stated in Section 2 Lemma 21 is proved using Lemma 20, also below.
@@
Assume that $G$ is $3$-connected and $3$-flattenable and contains a $3$-connected proper spanning subgraph. In [5] it was shown that any $3$-flattenable graph is a subgraph of a graph that can be constructed from $K_{4}$’s, $V_{8}$’s, and $C_{5}\times C_{2}$’s using $1$-sums, $2$-sums, and $3$-sums. Furthermore, since neither $V_{8}$ nor $C_{5}\times C_{2}$ contains a $K_{3}$ subgraph, these graphs can only be involved in $1$-sums and $2$-sums. Hence, $G$ is a subgraph of some graph $H$ that can be constructed as described. Furthermore, since $G$ is $3$-connected, $H$ is either a $3$-tree, $V_{8}$, or $C_{5}\times C_{2}$. In the latter two cases, every vertex of $H$ has degree three. However, this implies that no proper spanning subgraph of $G$ is $3$-connected, which contradicts our assumption. Thus, $H$ is a $3$-tree, and so the lemma is proved. ∎
Lemma 21 (Edge additions to links that preserve partial $3$-tree).

### v2

@@ 4.1   Proofs of Tool Lemmas 2 and 3
The lemma is immediate if $xy$ is an edge of $G\cup f$. Otherwise, assume to the contrary that $f$ is preserved in some forbidden minor $[G\cup\{f,xy\}]$. Let $[G\cup f]_{1}$ be an $f$-winged graph minor such that $xy$ is one of its pins. By Lemma 6, we can assume wlog that $[G\cup f]_{1}$ has wing tips $s$ and $t$. Since $[G\cup f]_{1}$ has at least two pins, by Proposition 3, $C=\{x,y,t\}$ is a separator of $G\cup\{f,xy\}$. Furthermore, since every forbidden minor is $4$-connected and has at least five vertices, there is a unique $C$-component $H$ of $G\cup\{f,xy\}$ that contains some vertex $w$ such that $[w]$ is not contained in $[C]$. If $H$ contains $f$, then the paths in $G\cup f$ given by the existence of $[G\cup f]_{1}$ imply the existence of some minor $[G\cup f]_{2}$ in which $H$ is preserved and $[C]_{2}$ is a clique. Combining these facts shows that $[G\cup f]_{2}$ has an $[f]_{2}$-preserving forbidden minor, and thus $G\cup f$ has an $f$-preserving forbidden minor, which is a contradiction.
Next, if $H$ does not contain $f$, then the existence of $[G\cup\{f,xy\}]$ implies the existence of a minor $[G\cup\{f,xy\}]_{4}$ that preserves $H$, has an $[f]_{4}$-preserving forbidden minor, and such that, for every vertex $w\in V(G)\setminus V(H)$, $[w]_{4}$ is contained in $[C]_{4}$. Hence, both endpoints of $[f]_{4}$ are contained $[C]_{4}$. Since $s$ and $t$ are wing tips of $[G\cup f]_{1}$, it cannot be that $[f]_{4}=[xy]_{4}$. Hence, assume wlog that $[f]_{4}=[xt]_{4}$. Then, we must have $[s]_{4}=[x]_{4}$. Since $C$ is a separator of $G\cup f$, $[y]_{1}^{-1}\setminus V(H)$ must be a subset of $[x]_{4}^{-1}\cup[y]_{4}^{-1}$. Therefore, it is easy to see that $xy$ is doubled in $[G\cup\{f,xy\}]_{4}$, and so $G\cup f$ clearly has an $f$-preserving forbidden minor, which is a contradiction. Thus, $G\cup\{f,xy\}$ has no $f$-preserving forbidden minor. ∎
@@
Consider any winged graph pair $(G,f)$ and any pin $xy$ of an $f$-winged graph minor $[G\cup f]$. If the converse direction of Theorem 4 is true for any graph-nonedge pair whose graph has strictly fewer vertices than $G$, then $(G\setminus\{[f]^{-1},xy\},xy)$ has the $3$-SIP.
@@
Note that $\{x,y\}$ is a clique separator of $(G\setminus[f]^{-1})\cup xy$. Hence, by Lemma 3, it suffices to show that show that $(H\setminus xy,xy)$ has the $3$-SIP for each $\{x,y\}$-component $H$ of $(G\setminus[f]^{-1})\cup xy$ that contains $xy$. If $(H\setminus xy,xy)$ does not have the $3$-SIP, then, since $|V(H)|<|V(G)|$, the contrapositive of the converse direction of Theorem 4 shows that $H$ has an $xy$-preserving forbidden minor. However, this fact along with the paths in $G\setminus[f]^{-1}$ given by the existence of $[G\cup f]$ imply that $G\cup f$ has an $f$-preserving forbidden minor, which is a contradiction. Thus, $(H\setminus xy,xy)$ has the $3$-SIP, and the proof is complete. ∎
Now we are ready to state and prove Proposition 4.
Proposition 4 (Chained winged-graph minors).
@@
Let $(G^{\prime},f)$ be any $f$-winged graph pair with an $f$-winged graph minor $[G^{\prime}\cup f]$ such that (i) $G$ is a spanning subgraph of $G^{\prime}$ and $(G,f)$ has the $3$-SIP if $(G^{\prime}\setminus[f]^{-1},[f]^{-1})$ is $3$-Cayley-connected and (ii) the number $k$ of pins of $[G^{\prime}\cup f]$ that are edges of $G^{\prime}\cup f$ is maximized. Note that $[G^{\prime}\cup f]$ has at least two pins, by Proposition 3. If $[G^{\prime}\cup f]$ is chained, then we are done. Otherwise, some pin $xy$ of $[G^{\prime}\cup f]$ is a nonedge of $G^{\prime}\cup f$. Using Lemma 10, we see that $(G^{\prime}\cup xy,f)$ is a winged graph pair. Also, the minor $[G^{\prime}\cup\{f,xy\}]_{1}$ such that, for any two vertices $u,v\in V(G^{\prime})$, $[u]_{1}=[v]_{1}$ if and only if $[u]=[v]$ is an $f$-winged graph minor. Note that $G$ is a spanning subgraph of $G^{\prime}\cup xy$ and the number of pins of $[G^{\prime}\cup\{f,xy\}]_{1}$ that are edges of $G^{\prime}\cup\{f,xy\}$ is strictly greater than $k$. Furthermore, Lemma 11 shows that $(G^{\prime}\setminus[f]^{-1},xy)$ has the $3$-SIP. Therefore, if $((G^{\prime}\setminus[f]^{-1})\cup xy,[f]^{-1})$ is $3$-Cayley connected, then Lemma 2 shows that $(G^{\prime}\setminus[f]^{-1},[f]^{-1})$ is $3$-Cayley connected. Consequently, $(G,f)$ has the $3$-SIP if $((G^{\prime}\setminus[f]^{-1})\cup xy,[f]^{-1})$ is $3$-Cayley connected. Combining the above facts contradicts the maximality of $(G^{\prime},f)$. Thus, every pin of $[G^{\prime}\cup f]$ is an edge of $G^{\prime}\cup f$, and the proof is complete. ∎
Now that a special structure has been established for $G_{\star}\cup f_{\star}$ in the first part of the proof of the converse direction of the main Theorem 4, we move to the second part of the proof. Despite this special structure, almost any linkage of $G_{\star}$ has a disconnected CS. Hence the main work of the second part of the proof is to establish that each connected component of the CS attains the same interval of lengths for $f_{\star}$.
For any pin $xy$ of the chained $f_{\star}$-winged graph minor $[G_{\star}\cup f_{\star}]$, there exist unique $\{x,y\}$-components $H_{u}(xy)$ and $H_{v}(xy)$ of $G_{\star}\setminus[f_{\star}]^{-1}$ that contain $u$ and $v$, respectively. Since every pin of $[G_{\star}\cup f_{\star}]$ is an edge of $G_{\star}\cup f_{\star}$, each pin of $[G_{\star}\cup f_{\star}]$ other than $xy$ is contained in either $H_{u}(xy)$ or $H_{v}(xy)$, but not both. Hence, we define the ordered sets of pins $\{x_{i}y_{i}\}$ of $[G_{\star}\cup f_{\star}]$ to be the set of all pins of $[G_{\star}\cup f_{\star}]$ indexed such that $H_{v}(x_{i}y_{i})$ contains $x_{i+1}y_{i+1}$. The links of $[G_{\star}\cup f_{\star}]$ are the graphs $H_{u}(x_{1}y_{1})$ and $H_{v}(x_{n}y_{n})$ and, for each $i<n$, the $\{x_{i+1},y_{i+1}\}$-component of $H_{v}(x_{i}y_{i})$ that contains $x_{i}y_{i}$. Proposition 5 shows that $(G_{\star},\{ux_{2},uy_{2}\})$ is $3$-Cayley-connected. See Figure 12(d). This proposition establishes that although the link of $G_{\star}$ between any consecutive pins $x_{i-1}y_{i-1}$ and $x_{i}y_{i}$ could force disconnectedness of the overall CS, the connected components of the CS can be viewed as partial reflections each of which attains the same connected set of lengths for the pair of nonedges $\{ux_{i},uy_{i}\}$. The same statement holds for the pair of nonedges $\{vx_{i},vy_{i}\}$ for any $i$, resulting in the same interval of lengths for $f_{\star}$.
Proposition 5 is proved using Lemmas 12 and 13, stated below and proved in A.4.
@@
Recall the $d$-covering map property defined in Section 2.
Lemma 13 is proved using the tool Theorems 5 and 6, which were stated in the Contributions Section 2 as independently interesting byproducts of the main result Theorem 4. We devote the next subsection to their proofs before stating and proving Proposition 5 and completing the proof of the converse direction of the main Theorem 4.
4.2   Proofs of Tool Theorems 5 and 6 concerning partial 3-trees and covering maps
The proof of Theorem 5 requires Lemmas 14 and 15, below, which may themselves be independently interesting results about partial $3$-trees.
@@ 4.2   Proofs of Tool Theorems 5 and 6 concerning partial 3-trees and covering maps
Assume that $G$ is $3$-connected and $3$-flattenable and contains a $3$-connected proper spanning subgraph. In [17] it was shown that a graph is $3$-flattenable it and only if it is a subgraph of a graph constructed from $K_{4}$s, $V_{8}$s, and $C_{5}\times C_{2}$s via $1$-sums, $2$-sums, and $3$-sums. Since neither $V_{8}$ nor $C_{5}\times C_{2}$ contains a $K_{3}$ subgraph, these graphs can only be involved in $1$-sums and $2$-sums. Hence, $G$ is a subgraph of some graph $H$ constructed as described above. Since $G$ is $3$-connected, $H$ is either a $3$-tree, $V_{8}$, or $C_{5}\times C_{2}$. Recall from Section 2 that the latter two graphs are forbidden minors for partial $3$-trees. Therefore, $G$ is a partial $3$-tree unless it is once of these two graphs. However, observe that no proper spanning subgraph of these two graphs is $3$-connected, and so $G$ is neither of these graphs. This completes the proof. ∎
Lemma 15 (If $G\cup\{wu_{i}\}$ with |$|\{wu_{i}\}|\in\{2,3\}$ has a forbidden minor, then so does $G\cup wu_{i}$ for some $wu_{i}$).
Let $G$ be a $3$-connected graph and $F=\{wu_{i}\}$ be some size two or three set of its nonedges. If there exists a forbidden minor $[G\cup F]$ in which $[w]$ and each $[u_{i}]$ are pairwise distinct vertices and each $wu_{i}$ is retained, then $G\cup wu_{i}$ has a forbidden minor for some $wu_{i}$.
The proof of Lemma 15 is given in A.2.
Proof of Theorem 5.
Let $G$ be a $3$-connected partial $3$-tree with a nonedge set $\{wu_{i}\}$ for which $G\cup wu_{i}$ is a partial $3$-tree for each $wu_{i}$. If $G\cup\{wu_{i}\}$ is $3$-flattenable, then the theorem follows from Lemma 14. Otherwise, let $F\subseteq\{wu_{i}\}$ be a minimum-sized set such that there exists a forbidden minor $[G\cup F]$. By our assumptions, we have $|F|\geq 2$. We first reduce the problem to the case where some nonedge in $F$ is retained in $[G\cup F]$. If some $f\in F$ is contracted in every forbidden minor of $G\cup F$, then Proposition 1 shows that $(G\cup(F\setminus\{f\}),f)$ is a winged graph pair with an $f$-winged graph minor. Hence, either Proposition 3 or Lemma 6 shows that $G\cup(F\setminus\{f\})$ has a separator of size two, a contradiction. Therefore, every nonedge in $F$ is preserved in some forbidden minor of $G\cup F$. This allows us to assume wlog that the set $X$ of all nonedges in $F$ that are preserved in $[G\cup F]$ is non-empty. The minimality condition on $F$ implies that each nonedge in $X$ is retained in $[G\cup F]$ and any two of these nonedges $wu_{i}$ and $wu_{j}$ are such that $[u_{i}]\neq[u_{j}]$.
Next, since $G$ is connected, each vertex of any forbidden minor has degree four, and $w$ is an endpoint of every nonedge in $F$, we have $|X|\leq 3$. If $|X|>1$, then Lemma 15 can be used to show the existence of the proper subset $F^{\prime}\subsetneq F$ such that $G\cup F^{\prime}$ has a forbidden minor, a contradiction. Therefore, wlog we have $X=\{wu_{1}\}$. Let $H$ be the subgraph of $G$ induced by $[w]^{-1}$ and let $H_{w}$ be its connected component that contains $w$. Also, let $\{[v_{i}]\}$ be the neighborhood of $[w]$ in $[G\cup F]\setminus[u_{1}]$. There are several cases.
If each set $[v_{i}]^{-1}$ contains some vertex whose neighborhood in $G$ contains some vertex in $H_{w}$, then $G\cup wu_{1}$ clearly has a forbidden minor, a contradiction. Otherwise, let $\{J_{i}\}$ be the set of connected components of $H\setminus H_{w}$. The minimality of $F$ implies that each $J_{i}$ contains both exactly one vertex $u_{j}$ such that $wu_{j}\in F$ and some vertex whose neighborhood in $G$ contains some vertex in at least one of the sets $[v_{i}]^{-1}$ or $[u_{1}]^{-1}$. If exactly two sets in $\{[v_{i}]^{-1}\}$, say $[v_{1}]^{-1}$ and $[v_{2}]^{-1}$, each contain some vertex that neighbora some vertex in $H_{w}$, then wlog $J_{1}$ contains some vertex whose neighborhood in $G$ contains some vertex in $[v_{3}]^{-1}$. The minor $[G\cup F]_{1}$ obtained via the component exchange from $[w]^{-1}$ to $[v_{3}]^{-1}$ that exchanges $J_{1}$ either is or has a forbidden minor in which at least two pairs in $F$ are retained, which is a previously handled case.
Consequently, assume that exactly one set in $\{[v_{i}]^{-1}\}$, say $[v_{1}]^{-1}$, contains some vertex whose neighborhood in $G$ contains some vertex in $H_{w}$. Then, $[v_{2}]^{-1}$ and $[v_{3}]^{-1}$ each contain some vertex that neighbors some vertex in some $J_{i}$. Assume the set $\{J_{i}\}$ contains more than one component. If only one component, say $J_{1}$, contains vertices with neighbors in $[v_{2}]^{-1}\cup[v_{3}]^{-1}$, then we arrive at a similar contradiction by considering the minor obtained via the component exchange from $[w]^{-1}$ to either $[u_{1}]^{-1}$ or $[v_{3}]^{-1}$, whichever contains a vertex with a neighbor in $J_{2}$, that exchanges $J_{2}$. Otherwise, wlog $J_{1}$ contains some vertex that neighbors some vertex in $[v_{2}]^{-1}$ and $J_{2}$ contains some vertex that neighbors some vertex in $[v_{3}]^{-1}$, and we arrive at a similar contradiction by considering the minor obtained via the component exchange from $[w]^{-1}$ to $[v_{3}]^{-1}$ that exchanges $J_{2}$.
Lastly, assume $J_{1}$ is the only component in $\{J_{i}\}$. First, we argue that we can assume $w$ is the only vertex in $H_{w}$. Since $G$ is $3$-connected, each connected component $H^{\prime}$ of $H_{w}\setminus w$ must contain some vertex that neighbors some vertex outside of $H$. Our above assumptions imply that this vertex outside $H$ is contained in $[v_{1}]^{-1}$. Hence, we can instead work with the minor obtained from $[G\cup F]$ via the component exchange from $[w]^{-1}$ to $[v_{1}]^{-1}$ that exchanges $H^{\prime}$. Repeating this argument for the new minor allows us to make the desired assumption on $H_{w}$.
Next, since $G$ is $3$-connected, there exist three vertex-disjoint paths $P_{a}$, $P_{b}$, and $P_{c}$ in $G$ between $w$ and $u_{1}$, where $a$, $b$, and $c$ are the first vertices along these paths, respectively, not in $[w]^{-1}$. Let $K$ be the subgraph of $G$ induced by $[v_{1}]^{-1}$. By our above assumptions, $K$ contains $a$, $b$, and $c$. Let $a^{\prime}$ be the first vertex in $P_{a}$ after $a$ that is not contained in $K$, and let $R_{a}$ be the path in $P_{a}$ between $a$ and the vertex immediately preceding $a^{\prime}$. Define $R_{b}$ and $R_{c}$ similarly. Wlog, we can assume that some connected component $K_{bc}$ of $K\setminus R_{a}$ contains both $R_{b}$ and $R_{c}$. Let $V_{a}$ be the set of vertices of $G\setminus([w]^{-1}\cup[v_{1}]^{-1})$ that are adjacent to some vertex in $K\setminus K_{bc}$ and let $V_{bc}$ be the set of vertices of $G\setminus([w]^{-1}\cup[v_{1}]^{-1})$ that are adjacent to some vertex in $K_{bc}$. There are several cases.
For any neighbor $[x]$ of $[v_{1}]$ other than $[w]$, if $V_{a}$ is a subset of $[x]^{-1}$, then the minor obtained via the component exchange from $[v_{1}]^{-1}$ to $[x]^{-1}$ that exchanges $K\setminus K_{bc}$ either demonstrates that $G\cup(F\setminus\{wu_{1}\})$ has a forbidden minor, which contradicts the minimality of $F$, or places us in the above handled case where exactly two sets in $\{[v_{i}]^{-1}\}$ each contain some vertex that neighbors some vertex in $H_{w}$. Otherwise, let $[x]$ be a neighbor of $[v_{1}]$ such that $V_{bc}\cap[x]^{-1}$ is non-empty. If no such neighbor $[x]$ exists, then some vertex in $K_{bc}$ neighbors some vertex in $J_{1}$. In this case, the minor obtained via the component exchange from $[v_{1}]^{-1}$ to $[w]^{-1}$ that exchanges $K_{bc}$ demonstrates that $G\cup(F\setminus\{wu_{2}\})$ has a forbidden minor, which contradicts the minimality of $F$. Therefore, assume $[x]$ exists and, if possible, chooseit so that $V_{a}\cap[x]^{-1}$ is empty. If $[G\cup F]$ is $K_{5}$, then the minor $[G\cup F]_{1}$ obtained via the component exchange from $[v_{1}]^{-1}$ to $[x]^{-1}$ that exchanges $K_{bc}$ either violates the minimality of $F$ or places us in the above handled case where exactly two sets in $\{[v_{i}]^{-1}\}$ each contain some vertex that neighbors some vertex in $H_{w}$. If $[G\cup F]$ is $K_{2,2,2}$, then there is one additional possibility for $[G\cup F]_{1}$: it has a $K_{5}$ minor in which $wu_{1}$ is preserved. However, in this case we could have instead considered $[G\cup F]_{1}$ to begin with. Thus, all cases have been checked, and so the theorem is proved. ∎
The proof of Theorem 6 requires Lemmas 16 and 17, below, which concern the following objects. Consider a graph $G$ with $F$ being a set of its nonedges such that $G\cup F$ is a partial $3$-tree, a squared edge-length map $\ell$, and the Cayley map $\phi_{F}$ from the CS $\mathcal{C}^{3}(G,\ell)$ to the CCS $\Omega^{3}_{F}(G,\ell)$. Let $Y$ be the relative interior of the CCS. For any realization $p$ in the CS and any $K_{4}$ subgraph $H$ of $T$, the orientation of the set of points to which $V(H)$ maps under $p$ is given by a determinant on these points - this determinant is 00 if and only if these points are coplanar, and either positive or negative otherwise. The orientation vector $\tau(p)=(\tau_{i}(p))_{i\in\{1,\dots,m\}}$ is such that $\tau_{i}(p)=0$ if the orientation determinant for $V(A_{i})$ is 00 under $p$, $\tau_{i}(p)=-1$ if this determinant is negative, and $\tau_{i}(p)=1$ if this determinant is positive. Let $\Delta$ be the set of all vectors $\delta=(\delta_{i})_{i\in\{1,\dots,m\}}$ with $\delta_{i}\in\{-1,1\}$. The $\delta$-oriented CS $\mathcal{C}^{3}_{\delta}(G,\ell)$ is the open subset of $\phi^{-1}(Y)$ containing every realization whose orientation vector is $\delta$.
Lemma 16 (Realizations in the CCS relative interior preimage).
Let $G$ be a graph with $F$ being a subset of its nonedges such that $G\cup F$ is a $3$-tree, $\ell$ be a squared edge-length map, $\phi_{F}$ be the Cayley map from the CS $\mathcal{C}^{3}(G,\ell)$ to the CCS $\Omega^{3}_{F}(G,\ell)$, and $Y$ be the the relative interior of the CCS. Then:
  1. (i)
    For each $\delta\in\Delta$, the image of the $\delta$-oriented CS under $\phi_{F}$ is $Y$.
  2. (ii)
    Each connected component of the CS contains some realization in a $\delta$-oriented CS for some $\delta\in\Delta$.